# Possible high superconductivity in LaNiO under high pressure through manifestation of a nearly-half-filled bilayer Hubbard model

### 24 Jan 2024 v3

@@ Abstract
Method.—First, we perform first-principles calculation to obtain the band structure of LaNiO using the QUANTUM ESPRESSO code Giannozzi et al. (2017). Perdew-Burke-Ernzerhof parametrization of the generalized gradient approximation (PBE-GGA) Perdew et al. (1996) and the scalar-relativistic version of the optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) taken from PseudoDojo van Setten et al. (2018) are used. As for the structural parameters, we adopt those given in Ref. Sun et al. (2023) for LaNiO under the pressure of GPa, namely, the experimentally determined lattice constants as well as the atomic positions determined by theoretical optimization. Since the orthorhombicity at GPa is quite small (, where are lattice constants for space group ), we adopt a body-centered tetragonal structure (, Fig. 1(a)) as in LaCuO, with the in-plane lattice constant determined as an average of the original ones, i.e., (see supplemental material SM for details). We take 100 Ry plane-wave cutoff energy, a 12 12 12 -mesh, and an energy width of 0.02 Ry for Gaussian smearing.
@@
Results and Discussions.—In Fig. 2(a), we show the eigenvalue of the linearized Eliashberg equation at eV as a function of the band filling , denoted as “original model”. corresponds to the stoichiometric composition of the actual material, and is varied assuming a rigid band. In Fig. 2(d), we show the superconducting gap function of the present model at in the band representation. It can be seen that the gap function is large at portions of the band where the orbital component is large, and the bonding and antibonding portions of the bands (see Fig. 2(b)) have opposite signs of the gap.
@@ Acknowledgements.
We are supported by JSPS KAKENHI Grant No. JP22K03512 (H. S.) and JP22K04907 (K. K.). The computing resource is supported by the supercomputer system HOKUSAI in RIKEN, and the supercomputer system (system-B) in the Institute for Solid State Physics, the University of Tokyo.References
- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, Nature 518, 179 (2015).
- Hosono and Kuroki (2015) H. Hosono and K. Kuroki, Physica C: Superconductivity and its Applications 514, 399 (2015), superconducting Materials: Conventional, Unconventional and Undetermined.
- Arita et al. (2000) R. Arita, K. Kuroki, and H. Aoki, Journal of the Physical Society of Japan 69, 1181 (2000).
@@
- Mishra et al. (2016) V. Mishra, D. J. Scalapino, and T. A. Maier, Scientific Reports 6, 32078 (2016).
@@
- Matsumoto et al. (2020) K. Matsumoto, D. Ogura, and K. Kuroki, Journal of the Physical Society of Japan 89, 044709 (2020).
@@
- Wang et al. (2011) F. Wang, F. Yang, M. Gao, Z.-Y. Lu, T. Xiang, and D.-H. Lee, Europhysics Letters 93, 57003 (2011).
@@
- Hirschfeld et al. (2011) P. J. Hirschfeld, M. M. Korshunov, and I. I. Mazin, Reports on Progress in Physics 74, 124508 (2011).
@@
- Bang (2019) Y. Bang, Scientific Reports 9, 3907 (2019).
- Charnukha et al. (2015) A. Charnukha, D. V. Evtushinsky, C. E. Matt, N. Xu, M. Shi, B. Büchner, N. D. Zhigadlo, B. Batlogg, and S. V. Borisenko, Scientific Reports 5, 18273 (2015).
- Miao et al. (2015) H. Miao, T. Qian, X. Shi, P. Richard, T. K. Kim, M. Hoesch, L. Y. Xing, X.-C. Wang, C.-Q. Jin, J.-P. Hu, and H. Ding, Nature Communications 6, 6056 (2015).
@@
- Ogura (2019) D. Ogura, Theoretical Study of Electron Correlation Driven Superconductivity in Systems with Coexisting Wide and Narrow Bands, Doctor theis, Osaka Univ., Toyonaka, Osaka (2019).
@@
- Aoki (2020) H. Aoki, Journal of Superconductivity and Novel Magnetism 33, 2341 (2020).
- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, J. Cheng, D.-X. Yao, G.-M. Zhang, and M. Wang, Nature 621, 493 (2023).
@@
- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Computer Physics Communications 226, 39 (2018).
- (50) See Supplemental Material at https://github.com/hskkbr/la3ni2o7paper for details, which includes Refs. Wang et al. (2023); Geisler et al. (2023); LaBollita et al. (2023); Qu et al. (2023a); Jiang et al. (2023); Qu et al. (2023b); Lu et al. (2023); Oh and Zhang (2023); Kakoi et al. (2023).
@@
- Nakamura et al. (2021) K. Nakamura, Y. Yoshimoto, Y. Nomura, T. Tadano, M. Kawamura, T. Kosugi, K. Yoshimi, T. Misawa, and Y. Motoyama, Computer Physics Communications 261, 107781 (2021).
@@
- Nakamura et al. (2009) K. Nakamura, Y. Yoshimoto, T. Kosugi, R. Arita, and M. Imada, Journal of the Physical Society of Japan 78, 083710 (2009).
- Nakamura et al. (2008) K. Nakamura, R. Arita, and M. Imada, Journal of the Physical Society of Japan 77, 093711 (2008).
@@
- Fujiwara et al. (2003) T. Fujiwara, S. Yamamoto, and Y. Ishii, Journal of the Physical Society of Japan 72, 777 (2003).
@@
- Sakakibara et al. (2013) H. Sakakibara, K. Suzuki, H. Usui, K. Kuroki, R. Arita, D. J. Scalapino, and H. Aoki, Journal of Physics: Conference Series 454, 012021 (2013).
@@
- Momma and Izumi (2011) K. Momma and F. Izumi, Journal of Applied Crystallography 44, 1272 (2011).
@@
- Li et al. (2019) D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature 572, 624 (2019).
@@
- Qu et al. (2023a) X.-Z. Qu, D.-W. Qu, W. Li, and G. Su, (2023a), arXiv:2311.12769.
- Jiang et al. (2023) K. Jiang, Z. Wang, and F.-C. Zhang, (2023), arXiv:2308.06771.
- Qu et al. (2023b) X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, (2023b), arXiv:2307.16873.
- Lu et al. (2023) C. Lu, Z. Pan, F. Yang, and C. Wu, (2023), arXiv:2307.14965.
- Oh and Zhang (2023) H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).

### 06 Mar 2024 v5

@@ Abstract
Method.—First, we perform first-principles calculation to obtain the band structure of LaNiO using the QUANTUM ESPRESSO code Giannozzi et al. (2017). The Perdew-Burke-Ernzerhof parametrization of the generalized gradient approximation (PBE-GGA) Perdew et al. (1996) and the scalar-relativistic version of the optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) taken from PseudoDojo van Setten et al. (2018) are used. As for the structural parameters, we adopt those given in the extended data table 1 of Ref. Sun et al. (2023) for LaNiO under the pressure of GPa, namely, the experimentally determined lattice constants as well as the atomic positions determined by theoretical optimization. Since the orthorhombicity at GPa is quite small (, where are lattice constants for space group ), we adopt a body-centered tetragonal structure (, Fig. 1(a)) as in LaCuO, with the in-plane lattice constant determined as an average of the original ones, i.e., (see supplemental material SM for details, which includes Refs. Qu et al. (2023); Jiang et al. (2024); Qu et al. (2024); Lu et al. (2023); Oh and Zhang (2023); Kakoi et al. (2023)). Note that if optimizing not only the internal coordinates but also the lattice parameters within PBE-GGA, the structure indeed goes to Geisler et al. (2023); LaBollita et al. (2023), in agreement with a recent experiment Wang et al. (2023). We take 100 Ry plane-wave cutoff energy, a 12 12 12 -mesh, and an energy width of 0.02 Ry for Gaussian smearing.
@@
Results and Discussions.—In Fig. 2(a), we show the eigenvalue of the linearized Eliashberg equation at eV as a function of the band filling , denoted as “original model”. corresponds to the stoichiometric composition of the actual material, and is varied assuming a rigid band. In Fig. 2(d), we show the superconducting gap function of the present model at in the band representation. It can be seen that the gap function is large at portions of the band where the orbital component is large, and the bonding and antibonding portions of the bands [see Fig. 2(b)] have opposite signs of the gap.
@@ Acknowledgements.
We are supported by JSPS KAKENHI Grant No. JP22K03512 (H. S.) and JP22K04907 (K. K.). The computing resource is supported by the supercomputer system HOKUSAI in RIKEN, the supercomputer system (system-B) in the Institute for Solid State Physics, the University of Tokyo, and the supercomputer of Academic Center for Computing and Media Studies (ACCMS), Kyoto University.References
- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, Nature(London) 518, 179 (2015).
- Hosono and Kuroki (2015) H. Hosono and K. Kuroki, Physica(Amsterdam) 514C, 399 (2015), superconducting Materials: Conventional, Unconventional and Undetermined.
- Arita et al. (2000) R. Arita, K. Kuroki, and H. Aoki, J. Phys. Soc. Jpn. 69, 1181 (2000).
@@
- Mishra et al. (2016) V. Mishra, D. J. Scalapino, and T. A. Maier, Sci. Rep. 6, 32078 (2016).
@@
- Matsumoto et al. (2020) K. Matsumoto, D. Ogura, and K. Kuroki, J. Phys. Soc. Jpn. 89, 044709 (2020).
@@
- Wang et al. (2011) F. Wang, F. Yang, M. Gao, Z.-Y. Lu, T. Xiang, and D.-H. Lee, Europhys. Lett. 93, 57003 (2011).
@@
- Hirschfeld et al. (2011) P. J. Hirschfeld, M. M. Korshunov, and I. I. Mazin, Rep. Prog. Phys. 74, 124508 (2011).
@@
- Bang (2019) Y. Bang, Sci. Rep. 9, 3907 (2019).
- Charnukha et al. (2015) A. Charnukha, D. V. Evtushinsky, C. E. Matt, N. Xu, M. Shi, B. Büchner, N. D. Zhigadlo, B. Batlogg, and S. V. Borisenko, Sci. Rep. 5, 18273 (2015).
- Miao et al. (2015) H. Miao, T. Qian, X. Shi, P. Richard, T. K. Kim, M. Hoesch, L. Y. Xing, X.-C. Wang, C.-Q. Jin, J.-P. Hu, and H. Ding, Nat. Commun. 6, 6056 (2015).
@@
- Ogura (2019) D. Ogura, Theoretical Study of Electron Correlation Driven Superconductivity in Systems with Coexisting Wide and Narrow Bands, Doctoral theis, Osaka University, Toyonaka, Osaka (2019).
@@
- Aoki (2020) H. Aoki, J. Supercond. Novel Magn. 33, 2341 (2020).
- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, J. Cheng, D.-X. Yao, G.-M. Zhang, and M. Wang, Nature(London) 621, 493 (2023).
@@
- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Comput. Phys. Commun. 226, 39 (2018).
- Momma and Izumi (2011) K. Momma and F. Izumi, J. Appl. Crystallogr. 44, 1272 (2011).
- (51) See Supplemental Material at https://github.com/hskkbr/la3ni2o7paper for details, which includes Refs. Wang et al. (2023); Geisler et al. (2023); LaBollita et al. (2023); Qu et al. (2023); Jiang et al. (2024); Qu et al. (2024); Lu et al. (2023); Oh and Zhang (2023); Kakoi et al. (2023).
- Qu et al. (2023) X.-Z. Qu, D.-W. Qu, W. Li, and G. Su, (2023), arXiv:2311.12769.
- Jiang et al. (2024) K. Jiang, Z. Wang, and F.-C. Zhang, Chin. Phys. Lett. 41, 017402 (2024).
- Qu et al. (2024) X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, Phys. Rev. Lett. 132, 036502 (2024).
- Lu et al. (2023) C. Lu, Z. Pan, F. Yang, and C. Wu, (2023), arXiv:2307.14965.
- Oh and Zhang (2023) H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).
@@
- Nakamura et al. (2021) K. Nakamura, Y. Yoshimoto, Y. Nomura, T. Tadano, M. Kawamura, T. Kosugi, K. Yoshimi, T. Misawa, and Y. Motoyama, Comput. Phys. Commun. 261, 107781 (2021).
@@
- Nakamura et al. (2009) K. Nakamura, Y. Yoshimoto, T. Kosugi, R. Arita, and M. Imada, J. Phys. Soc. Jpn. 78, 083710 (2009).
- Nakamura et al. (2008) K. Nakamura, R. Arita, and M. Imada, J. Phys. Soc. Jpn. 77, 093711 (2008).
@@
- Fujiwara et al. (2003) T. Fujiwara, S. Yamamoto, and Y. Ishii, J. Phys. Soc. Jpn. 72, 777 (2003).
@@
- Sakakibara et al. (2013) H. Sakakibara, K. Suzuki, H. Usui, K. Kuroki, R. Arita, D. J. Scalapino, and H. Aoki, J. Phys. Conf. Ser. 454, 012021 (2013).
@@
- Li et al. (2019) D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature(London) 572, 624 (2019).